# Existence and large radial solutions for an elliptic system under finite new Keller-Osserman integral conditions

### v1

@@ 4 The set of central values: Nonemptiness and closedness
Uniform bounds on compact intervals. Fix $m\in\mathbb{N}$. Consider the scalar blow-up comparison problem (cf. (11) in Section 3):
$\begin{cases}z_{1}^{\prime\prime}(r)+\frac{n-1}{r}z_{1}^{\prime}(r)=p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g\left(f\left(z_{1}\left(r\right)\right)\right),\\ z_{1}(0)=c>0,\quad z_{1}^{\prime}(0)=0,\end{cases}$
and similarly for $z_{2}$ with $q$ and $f(g(\cdot))$. By the Keller-Osserman-type conditions and the integrability of $p,q$, there exists $R_{m}>m$ such that $z_{1m}$ and $z_{2m}$ (the blow-up solutions with suitably large initial data) are finite on $[0,m]$ and blow up at $r=R_{m}$.
By the monotonicity of $u_{k}$ and $v_{k}$ in $r$, and the comparison principle for the radial ODEs, we have
$u_{k}(r)\leq z_{1m}(r),\quad v_{k}(r)\leq z_{2m}(r),\qquad\forall r\in[0,m],\ \forall k.$
@@
$C^{1}$-bounds. From the radial form of (1) defined above:
$u_{k}^{\prime}\left(r\right)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)\,g(v_{k}(s))\,ds,$
and the uniform bound on $v_{k}$ together with assumption on $p$ yield a uniform bound on $u_{k}^{\prime}$ on $[0,m]$. Similarly for $v_{k}^{\prime}$. Hence $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded in $C^{1}([0,m])$.
Compactness and passage to the limit. By the Arzel-Ascoli theorem, for each $m$ there exists a subsequence (still denoted by $u_{k},v_{k}$) converging in $C^{1}([0,m])$ to some $(u^{(m)},v^{(m)})$. A diagonal extraction yields a pair $(u,v)$ such that
$(u_{k},v_{k})\rightarrow(u,v)\quad\text{in }C_{\mathrm{loc}}^{1}([0,\infty)).$
Passing to the limit in the integral formulation of (1) (valid for radial $C^{2}$-solutions), we find that $(u,v)$ is an entire positive radial solution with
$u(0)=a_{0},\quad v(0)=b_{0}.$
@@ 5 Largeness on the edge
Let $(a,b)\in E$ and $(u,v)$ be an entire solution with these central values. From Theorem 1.2 we have $\partial S\subset S$. Choose a sequence $(a_{k},b_{k})\notin S$ with $(a_{k},b_{k})\rightarrow(a,b)$ and $a_{k},b_{k}>0$. By Lemma 2.1, for each $k$ there is a maximal radius $R_{k}<\infty$ for the solution $(u_{k},v_{k})$ with $u_{k}(0)=a_{k}$, $v_{k}(0)=b_{k}$. By Lemma 2.2,
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$
If $\{R_{k}\}$ were bounded, a subsequence would converge to some $R<\infty$, and the local uniform convergence $(u_{k},v_{k})\rightarrow(u,v)$ would imply $(u,v)$ blows up at $r=R$, contradicting that $(u,v)$ is entire. Hence $R_{k}\rightarrow\infty$.
Fix $r>0$; for $k$ large enough we have $r<R_{k}$ and, by monotonicity in the initial data, $u_{k}(r)\leq u(r)$ and $v_{k}(r)\leq v(r)$. By Lemma 2.3 we have
$\Phi^{\prime\prime}(u_{k}(r))+\frac{n-1}{r}\Phi^{\prime}(u_{k}(r))\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),$ (13)
@@
$\Psi^{\prime\prime}(v_{k}(r))+\frac{n-1}{r}\Psi^{\prime}(v_{k}(r))\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$ (14)
Equations (13)–(14) can be rewritten as
$\left\{\begin{array}[]{c}\left(r^{n-1}\Phi^{\prime}(u_{k}\left(r\right))\right)^{\prime}\geq-r^{n-1}p(r)g(\frac{b}{f(a)}+\mathcal{L}_{q}),\\ \left(r^{n-1}\Psi^{\prime}(v_{k}\left(r\right))\right)^{\prime}\geq-r^{n-1}q(r)f(\frac{a}{g(b)}+\mathcal{L}_{p}).\end{array}\right$.
Integrating over $[0,r]$ with $0<r<R_{k}$ gives
$\left\{\begin{array}[]{c}\Phi^{\prime}(u_{k}\left(r\right))\geq-r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})ds,\\ \Psi^{\prime}(v_{k}\left(r\right))\geq-r^{1-n}\int_{0}^{r}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})ds.\end{array}\right$.
Next, integrate over $r,R_{k}]$. Since
$u_{k}(t)\rightarrow\infty\ \text{and}\ v_{k}(t)\rightarrow\infty\quad\text{as }t\rightarrow R_{k}^{-},$
we have
$\Phi(u_{k}(t))\rightarrow 0,\quad\Psi(v_{k}(t))\rightarrow 0\quad\text{as }t\rightarrow R_{k}^{-}.$
@@
$\left\{\begin{array}[]{c}-\Phi(u_{k}\left(r\right))\geq-\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\\ -\Psi(v_{k}\left(r\right))\geq-\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\end{array}\right$.
or equivalently,
$\left\{\begin{array}[]{c}\Phi(u_{k}\left(r\right))\leq\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt,\\ \Psi(v_{k}\left(r\right))\leq\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt.\end{array}\right$.
Since $\Phi$ and $\Psi$ are strictly decreasing (see (8)), we can invert them to obtain
$\left\{\begin{array}[]{c}u_{k}\left(r\right)\geq\Phi^{-1}\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\right),\\ v_{k}\left(r\right)\geq\Psi^{-1}\left(\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\right).\end{array}\right$.
Letting $k\rightarrow\infty$ gives $R_{k}\rightarrow R$ and $(u_{k},v_{k})\rightarrow(u,v)$ locally uniformly, hence
$\left\{\begin{array}[]{c}u\left(r\right)\geq\Phi^{-1}\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(\frac{b}{f(a)}+\mathcal{L}_{q})dsdt\right),\\ v\left(r\right)\geq\Psi^{-1}\left(\int_{r}^{R}t^{1-n}\int_{0}^{t}s^{n-1}q(s)f(\frac{a}{g(b)}+\mathcal{L}_{p})dsdt\right).\end{array}\right$.
Finally, letting $r\rightarrow R$ and using
$\lim_{t\rightarrow 0^{+}}\Phi^{-1}(t)=\lim_{t\rightarrow 0^{+}}\Psi^{-1}(t)=\infty,$
we deduce
$\lim_{r\rightarrow R}u(r)=\lim_{r\rightarrow R}v(r)=\infty.$
Since $(u,v)$ is entire with $(a,b)\in E\subset S$, we must have $R=\infty$. Thus $u(r),v(r)\rightarrow\infty$ as $r\rightarrow\infty$, i.e., $(u,v)$ is an entire large solution. ∎
Remark 2 (Criticality).

### v2

@@ 4 The set of central values: Nonemptiness and closedness
Uniform bounds on compact intervals. Fix $m\in\mathbb{N}$ arbitrarily. Since $(a_{k},b_{k})\to(a_{0},b_{0})$, the sequence $\{(a_{k},b_{k})\}$ is bounded. Without loss of generality, assume $a_{k}\leq A$ and $b_{k}\leq B$ for all $k$ and some constants $A,B>0$.
For each $k$, the solution $(u_{k},v_{k})$ satisfies, by the proof of Theorem 1.2, a comparison with auxiliary solutions $(z_{1,k},z_{2,k})$ solving:
$\begin{cases}z_{1,k}^{\prime\prime}+\frac{n-1}{r}z_{1,k}^{\prime}=p(r)g(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q})\,g\left(f\left(z_{1,k}\right)\right),\\ z_{1,k}(0)=c_{k}>a_{k},\quad z_{1,k}^{\prime}(0)=0,\end{cases}$ (17)
and similarly for $z_{2,k}$. However, to obtain uniform bounds independent of $k$, we construct a single comparison solution that dominates all $(u_{k},v_{k})$ simultaneously.
Let $\tilde{a}=\sup_{k}a_{k}\leq A$ and $\tilde{b}=\sup_{k}b_{k}\leq B$ (both finite by boundedness). Choose constants $\bar{c}>\tilde{a}$ and $\bar{d}>\tilde{b}$. Consider the auxiliary scalar problems:
$\begin{cases}\bar{z}_{1}^{\prime\prime}+\frac{n-1}{r}\bar{z}_{1}^{\prime}=p(r)g(\frac{B}{f(\inf_{k}a_{k})}+\mathcal{L}_{q})\,g\left(f\left(\bar{z}_{1}\right)\right),\\ \bar{z}_{1}(0)=\bar{c},\quad\bar{z}_{1}^{\prime}(0)=0,\end{cases}$
and
$\begin{cases}\bar{z}_{2}^{\prime\prime}+\frac{n-1}{r}\bar{z}_{2}^{\prime}=q(r)f(\frac{A}{g(\inf_{k}b_{k})}+\mathcal{L}_{p})f\left(g\left(\bar{z}_{2}\right)\right),\\ \bar{z}_{2}(0)=\bar{d},\quad\bar{z}_{2}^{\prime}(0)=0.\end{cases}$
By the Keller-Osserman-type conditions (F3) and the decay assumptions (PQ), these problems admit entire positive radial solutions $\bar{z}_{1}$ and $\bar{z}_{2}$ (see [6, 2]).
By the comparison principle (as in the proof of Theorem 1.2), since $a_{k}<\bar{c}$ and $b_{k}<\bar{d}$ for all $k$, and the right-hand sides of the equations for $(u_{k},v_{k})$ are dominated by those for $(\bar{z}_{1},\bar{z}_{2})$ (due to monotonicity of $f,g$ and the choices of constants), we obtain
$u_{k}(r)\leq\bar{z}_{1}(r),\quad v_{k}(r)\leq\bar{z}_{2}(r),\qquad\forall r\geq 0,\ \forall k\in\mathbb{N}.$
In particular, for the fixed $m$, we have uniform bounds:
$u_{k}(r)\leq\bar{z}_{1}(m)=:M_{1}(m)<\infty,\quad v_{k}(r)\leq\bar{z}_{2}(m)=:M_{2}(m)<\infty,\quad\forall r\in[0,m],\ \forall k.$
@@
$C^{1}$-bounds. From the integral representation (6), we have
$u_{k}^{\prime}\left(r\right)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)\,g(v_{k}(s))\,ds,\quad r\in(0,m].$
Since $v_{k}(s)\leq M_{2}(m)$ for all $s\in[0,m]$ and $g$ is continuous and nondecreasing, we have
$g(v_{k}(s))\leq g(M_{2}(m))=:G_{m}<\infty,\quad\forall s\in[0,m],\ \forall k.$
Thus
$u_{k}^{\prime}(r)\leq r^{1-n}\int_{0}^{r}s^{n-1}p(s)G_{m}\,ds\leq G_{m}\mathcal{P}(m)<\infty,\quad\forall r\in[0,m],\ \forall k,$
where $\mathcal{P}(m)<\infty$ by assumption (PQ). By symmetry,
$v_{k}^{\prime}(r)\leq f(M_{1}(m))\mathcal{Q}(m)<\infty,\quad\forall r\in[0,m],\ \forall k.$
Hence $\{u_{k}\}$ and $\{v_{k}\}$ are uniformly bounded in $C^{1}([0,m])$.
Moreover, from the equations (5), for $r\in[\delta,m]$ with $\delta>0$,
$u_{k}^{\prime\prime}(r)=-\frac{n-1}{r}u_{k}^{\prime}(r)+p(r)g(v_{k}(r))\leq\frac{n-1}{\delta}G_{m}\mathcal{P}(m)+\|p\|_{L^{\infty}([0,m])}G_{m},$
so $\{u_{k}^{\prime\prime}\}$ is uniformly bounded on $\delta,m]$. Similarly for $v_{k}^{\prime\prime}$. Thus $\{(u_{k},v_{k})\}$ is equicontinuous in $C^{1}$ on $\delta,m]$ for any $\delta>0$.
Compactness and passage to the limit. By the Arzelà-Ascoli theorem, for each $m\in\mathbb{N}$ there exists a subsequence (denote it by $\{u_{k_{j}^{(m)}},v_{k_{j}^{(m)}}\}$) converging in $C^{1}([0,m])$ to some $(u^{(m)},v^{(m)})\in C^{1}([0,m])$. By a standard diagonal argument, we extract a subsequence $\{k_{\ell}\}_{\ell=1}^{\infty}$ (still denoted by $\{k\}$ for simplicity) such that
$(u_{k},v_{k})\rightarrow(u,v)\quad\text{in }C_{\mathrm{loc}}^{1}([0,\infty)),$
i.e., $(u_{k},v_{k})\to(u,v)$ in $C^{1}([0,m])$ for every $m\in\mathbb{N}$.
We now show that $(u,v)$ is a solution of (1). Passing to the limit in the integral equations (6)–(7), using the continuity of $f,g,p,q$ and the dominated convergence theorem (justified by the uniform bounds), we obtain for every $r\geq 0$:
$u(r)$ $=\lim_{k\to\infty}u_{k}(r)=\lim_{k\to\infty}\left[a_{k}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(v_{k}(s))\,ds\,dt\right]$
$=a_{0}+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)g(v(s))\,ds\,dt,$
and similarly for $v$. Thus $(u,v)$ satisfies (6)–(7), hence is a $C^{2}$ positive radial solution of (1) with
$u(0)=a_{0},\quad v(0)=b_{0},\quad u^{\prime}(0)=v^{\prime}(0)=0.$
@@ 5 Largeness on the edge
Let $(a,b)\in E=\partial S\cap(0,\infty)^{2}$ be a boundary point of $S$ in the positive quadrant, and let $(u,v)$ be an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$.
Since $S$ is closed (Theorem 1.3), we have $\partial S\subset S$, so $(a,b)\in S$ and the solution $(u,v)$ exists and is entire.
Step 1: Approximating sequence. Since $(a,b)\in\partial S$, there exists a sequence $(a_{k},b_{k})\notin S$ with $(a_{k},b_{k})\rightarrow(a,b)$ as $k\to\infty$, and $a_{k},b_{k}>0$ for all $k$.
For each $k$, by Lemma 2.1, there exists a maximal radius $R_{k}\in(0,\infty]$ and a positive radial solution $(u_{k},v_{k})\in C^{2}([0,R_{k}))^{2}$ of (1) with $u_{k}(0)=a_{k}$, $v_{k}(0)=b_{k}$, and $u_{k}^{\prime}(0)=v_{k}^{\prime}(0)=0$.
Since $(a_{k},b_{k})\notin S$, the solution $(u_{k},v_{k})$ cannot be extended to all of $0,\infty)$, hence $R_{k}<\infty$ for all $k$. By Lemma 2.2,
$\lim_{r\rightarrow R_{k}^{-}}u_{k}(r)=\lim_{r\rightarrow R_{k}^{-}}v_{k}(r)=\infty.$ (18)
Step 2: The sequence $\{R_{k}\}$ is unbounded. Suppose, for contradiction, that $\{R_{k}\}$ is bounded. Then there exists $R_{\star}<\infty$ such that $R_{k}\leq R_{\star}$ for all $k$. Passing to a subsequence if necessary, we may assume $R_{k}\to R_{0}\leq R_{\star}$ as $k\to\infty$.
[…]
But $(u,v)$ is the unique solution with these initial values (by uniqueness in Lemma 2.1), so $\tilde{u}=u$ and $\tilde{v}=v$ on $0,R_{0})$. Thus $u_{k}\to u$ and $v_{k}\to v$ locally uniformly on $0,R_{0})$.
However, from (18), $u_{k}(r)\to\infty$ as $r\to R_{k}^{-}\to R_{0}^{-}$, which would imply $u(r)\to\infty$ as $r\to R_{0}^{-}$, contradicting that $(u,v)$ is entire. Therefore, $\{R_{k}\}$ is unbounded, i.e., $R_{k}\rightarrow\infty$ as $k\to\infty$.
Step 3: Application of the Keller-Osserman transforms. Now that we have established $R_{k}\to\infty$, we apply the transforms $\Phi$ and $\Psi$ to derive lower bounds on $u$ and $v$.
Fix $r>0$ arbitrarily. For $k$ sufficiently large, we have $r<R_{k}$, so $(u_{k},v_{k})$ is well-defined on $[0,r]$.
By Lemma 2.3, specifically inequality (11), the functions $u_{k}$ and $v_{k}$ satisfy for all $\rho\in(0,R_{k})$:
$\Phi^{\prime\prime}(u_{k}(\rho))+\frac{n-1}{\rho}\Phi^{\prime}(u_{k}(\rho))\ \geq\ -p(\rho)\,g\!\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right),$ (19)
@@
$\Psi^{\prime\prime}(v_{k}(\rho))+\frac{n-1}{\rho}\Psi^{\prime}(v_{k}(\rho))\ \geq\ -q(\rho)\,f\!\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right).$ (20)
Since $(a_{k},b_{k})\to(a,b)$ and $f,g$ are continuous, the constants $\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\to\frac{b}{f(a)}+\mathcal{L}_{q}$ and $\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\to\frac{a}{g(b)}+\mathcal{L}_{p}$ as $k\to\infty$. For $k$ sufficiently large, these constants are bounded, say by $C_{1}$ and $C_{2}$ respectively.
Inequalities (19)–(20) can be rewritten in divergence form:
$\left\{\begin{array}[]{c}\left(\rho^{n-1}\Phi^{\prime}(u_{k}(\rho))\right)^{\prime}\geq-\rho^{n-1}p(\rho)g(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}),\\ \left(\rho^{n-1}\Psi^{\prime}(v_{k}(\rho))\right)^{\prime}\geq-\rho^{n-1}q(\rho)f(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}).\end{array}\right$. (21)
Step 4: Integration and derivation of lower bounds. Integrating (21) over $[0,r]$ with $0<r<R_{k}$, and using $\Phi^{\prime}(u_{k}(0))=-\frac{u_{k}^{\prime}(0)}{g(f(u_{k}(0)))}=0$ (since $u_{k}^{\prime}(0)=0$), we obtain
$r^{n-1}\Phi^{\prime}(u_{k}(r))\geq-\int_{0}^{r}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds.$ (22)
Dividing by $r^{n-1}$ yields
$\Phi^{\prime}(u_{k}(r))\geq-r^{1-n}\int_{0}^{r}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds.$
Similarly,
$\Psi^{\prime}(v_{k}(r))\geq-r^{1-n}\int_{0}^{r}s^{n-1}q(s)f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)ds.$
Now, we integrate over $r,R_{k}]$. Since $u_{k}(t)\rightarrow\infty$ and $v_{k}(t)\rightarrow\infty$ as $t\rightarrow R_{k}^{-}$ (by (18)), and since $\Phi(t)=\int_{t}^{\infty}\frac{ds}{g(f(s))}\to 0$ as $t\to\infty$ (by the finite integral condition (8)), we have
$\lim_{t\to R_{k}^{-}}\Phi(u_{k}(t))=0,\quad\lim_{t\to R_{k}^{-}}\Psi(v_{k}(t))=0.$ (23)
Integrating the inequality for $\Phi^{\prime}(u_{k}(\rho))$ over $\rho\in[r,R_{k})$, we obtain
$0-\Phi(u_{k}(r))$ $=\lim_{t\to R_{k}^{-}}\Phi(u_{k}(t))-\Phi(u_{k}(r))=\int_{r}^{R_{k}}\Phi^{\prime}(u_{k}(\rho))\,d\rho$
$\geq-\int_{r}^{R_{k}}\rho^{1-n}\int_{0}^{\rho}s^{n-1}p(s)g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)ds\,d\rho.$
@@
$\Phi(u_{k}(r))\leq g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt.$ (24)
Similarly,
$\Psi(v_{k}(r))\leq f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt.$ (25)
Step 5: Inversion and passage to the limit. Since $\Phi$ and $\Psi$ are strictly decreasing (see (9)), they are invertible. From (24)–(25), we obtain
$u_{k}(r)$ $\geq\Phi^{-1}\left(g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\right),$ (26)
$v_{k}(r)$ $\geq\Psi^{-1}\left(f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\right).$ (27)
Now, let $k\to\infty$. We have:
  • $R_{k}\to\infty$ (Step 2);
  • $(a_{k},b_{k})\to(a,b)$, so $g\left(\frac{b_{k}}{f(a_{k})}+\mathcal{L}_{q}\right)\to g\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)=:G_{0}$ and $f\left(\frac{a_{k}}{g(b_{k})}+\mathcal{L}_{p}\right)\to f\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)=:F_{0}$ by continuity;
  • For any fixed $r>0$,
    $\int_{r}^{R_{k}}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\to\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt=\mathcal{P}(\infty)-\mathcal{P}(r)$
    as $k\to\infty$, where $\mathcal{P}(\infty)=\mathcal{L}_{p}(n-2)<\infty$ by assumption (PQ). Similarly for $q$.
By the continuity of $\Phi^{-1}$ and passing to the limit in (26) as $k\to\infty$, we obtain (noting that $u_{k}\to u$ locally uniformly from Step 2’s argument):
$u(r)\geq\Phi^{-1}\left(G_{0}\left[\mathcal{L}_{p}(n-2)-\mathcal{P}(r)\right]\right)=\Phi^{-1}\left(G_{0}\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\right).$ (28)
Similarly,
$v(r)\geq\Psi^{-1}\left(F_{0}\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\right).$ (29)
Step 6: Blow-up at infinity. Since $p,q$ are not both identically zero and $\min(p,q)$ does not have compact support (assumption (PQ)), at least one of the integrals $\int_{0}^{\infty}s^{n-1}p(s)\,ds$ or $\int_{0}^{\infty}s^{n-1}q(s)\,ds$ is positive. Thus, as $r\to\infty$,
$\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,ds\,dt\to 0\quad\text{and/or}\quad\int_{r}^{\infty}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt\to 0.$
Since $\Phi(t)\to 0$ as $t\to\infty$ (by (8)), we have $\Phi^{-1}(s)\to\infty$ as $s\to 0^{+}$. Similarly for $\Psi^{-1}$. Therefore, from (28)–(29), as $r\to\infty$,
$u(r)\geq\Phi^{-1}(G_{0}\cdot[\text{positive term}\to 0])\to\infty,\quad v(r)\geq\Psi^{-1}(F_{0}\cdot[\text{positive term}\to 0])\to\infty.$
Hence,
$\lim_{r\to\infty}u(r)=\infty\quad\text{and}\quad\lim_{r\to\infty}v(r)=\infty,$
proving that $(u,v)$ is a large solution. This completes the proof. ∎
Remark 5.1 (Criticality).
@@
6 Compatibility between Theorem 1.4 and Theorem 2.4 of [3]
6.1 General framework
Both ELRS (Theorem 1.4) and Theorem 2.4 of [3] study radial elliptic systems of the form (1), where in [3] $S_{k_{i}}$ denotes the $k_{i}$-Hessian operator (the case $k_{i}=1$ being the Laplacian). Here $p,q$ are nonnegative radial weights and $f,g$ are positive nonlinearities.
  • In ELRS, finite Keller–Osserman-type conditions are imposed to ensure the existence of both bounded entire and large (blow-up at infinity) radial solutions, depending on the central values $(u(0),v(0))$.
  • In T2.4, the focus is on positive bounded entire radial solutions for the Hessian system with gradient term, under radial symmetry, a factorization structure for $f,g$, and a strict hierarchy between the functionals $\mathcal{P},\mathcal{Q},\Phi,\Psi$.
6.2 A common particular example
Let
$g(s)$ $=$ $s^{\alpha},\quad f(s)=s^{\beta},\quad\alpha=\beta=\frac{6}{5},\quad s\geq 0,$
$p(r)$ $=$ $q(r)=\frac{1}{(1+r)^{\gamma}},\quad\gamma=5,\quad r\geq 0.$
Then:
  • $f$ and $g$ are continuous, positive, strictly increasing on $(0,\infty)$;
  • the classical Keller–Osserman condition holds;
  • the potential integrals are finite since $p,q\sim r^{-5}$;
  • $\alpha\beta>1$ ensures the finiteness of the $\circ$-functionals;
  • the strict hierarchy $P_{i,j}(\infty)<H_{i,j}(\infty)$ required in T2.4 is satisfied.
6.3 Applicability of the two results
For ELRS:
The finite Keller–Osserman conditions together with finite potential integrals place this example in the class where there exist both bounded entire solutions (central values in the interior of $S$) and large solutions (central values on the boundary $E=\partial S$). ELRS distinguishes these regimes by the location of the central value.
For T2.4:
All hypotheses (radial symmetry, factorization, monotonicity, strict hierarchy) are met. The theorem yields a positive bounded entire radial solution for central values satisfying the strict inequality in the hypotheses. Such values lie in the interior of $S$ in the ELRS framework.
6.4 Why there is no contradiction
Let $S$ be the set of central values for which entire solutions exist, and $E=\partial S\cap(0,\infty)^{2}$ its “edge” in the positive quadrant.
  • T2.4 produces bounded entire solutions for $(u(0),v(0))$ in the interior of $S$.
  • ELRS asserts that for $(u(0),v(0))\in E$ any entire solution is large, while for interior points bounded entire solutions also exist.
In the above example, choosing central values according to T2.4 (strict inequality) places the solution in $\mathrm{int}(S)$, outside the scope of the “large solution” conclusion of ELRS. The two results are therefore compatible: they describe different behaviours (bounded vs. large) in disjoint regions of the central value space, even for the same coefficients.
Moreover, $E$ is not empty under the stated hypotheses: taking the boundary $\partial S$ in $\mathbb{R}^{2}$ and intersecting with $(0,\infty)^{2}$ yields a nonempty edge where the qualitative behaviour changes. Points of $E$ mark the transition between bounded entire solutions (interior) and large solutions (edge), making $E$ both topologically and analytically significant.